OpenStudy (vortish):

solve and check 13x^2-20x=0

OpenStudy (anonymous):

You can factor out an x, giving you x(13x - 20) = 0 Set them both equal to 0 and solve! To check, simply plug them into the original equation.

OpenStudy (vortish):

can you walk threw it with me pls

OpenStudy (anonymous):

So you see in both of the terms, there is at least an x? You can factor that out. So now you have x(13x-20) = 0 You can seperate these, and set them both equal to 0, since 0 times anything is equal to 0 so you have x = 0 and 13x -20 = 0 13x = 20 x = 20/13 So, x will equal either 0 or 13/20.

OpenStudy (vortish):

so i understand this Jonathan so in the end the answer would be 13/20 and i see how you did this thanks